Suppose you have $ 150 comma 000 in​ cash, and you decide to borrow another $ 36 comma 000 at a 7 % interest rate to invest in the stock market. You invest the entire $ 186 comma 000 in a portfolio J with a 16 % expected return and a 27 % volatility.

a. What is the expected return and volatility​ (standard deviation) of your​ investment?

b. What is your realized return if J goes up 19 % over the​ year?

c. What return do you realize if J falls by 14 % over the​ year?
Answer #1

1.
Expected return=186000/150000*16%-36000/150000*7%=18.1600%

Standard Deviation=186000/150000*27%=33.4800%

2.
=186000/150000*19%-36000/150000*7%=21.8800%

3.
=186000/150000*(-14%)-36000/150000*7%=-19.0400%